Three eggs were cracked. This was 1/6 of the total number of eggs in the carton.

A. How many eggs were in the carton?

B. What percent of the eggs in the carton were not cracked
1
Expert's answer
2020-05-04T18:24:58-0400

A. denote whole eggs as 1 and write a condition:

3 eggs - 1/6

x eggs - 1


now we can write the equation and solve it.

x*1/6 = 1*3

x = 3*6

x = 18 whole eggs that were in the carton


B. Now we know that total number of eggs are 18, 3 of them are cracked, so 18 - 3 = 15 number of not cracked eggs.

lets make a condition for it:

18 eggs - 100%

15 eggs - x


make an equation 18*x = 15*100

x = 1500/18

x = 83.33% of eggs that not cracked